OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 3.6x = 5.5 A. 15.3 B. 1.6 C. 1.5 D. 1.4 AND PLEASE EXPLAIN

OpenStudy (anonymous):

3.6x = 5.5 we want to isolate 'x', to do that we need to divide both sides by 3.6, can you do this for me? ^_^

OpenStudy (anonymous):

um ok so 3.6/3.6=1 and 5.5/3.6=1.52 right or do i round that so it equals 1.53 @DemolisionWolf

OpenStudy (anonymous):

x = 5.5/3.6 = 1.52777777777777777777 "round to the nearest tenth place" means we can only have one number after the decimal. so, 1.?

OpenStudy (anonymous):

1.5 then would be the answer right @DemolisionWolf